a pool party is occuring at the robertsons house this next holiday they want to invite 64 friends and family only 26 friends and 9 family members reply to the invite the robertsons are a family of 5 their pool is 35 feet long and is 10 feet deep when filled to capacity they are worried children under 5 will struggle in the pool so they only fill the pool to 5 feet the day of the party 53 people show up the son billy is 4.5 feet tall if billys height is half the width of the pool what is the volume of water in cubic feet the robinsons filled the pool with
1
Expert's answer
2020-01-28T05:44:03-0500
Answer  The length of the pool is =35 feet. 
        The depth of the pool is = 10 feet.
       Billy's height of the half of the width of pool 
Given  ,  (1/2)x =4.5    (x= width) 
             x= 4.5*2
           x= 9 feet    
so, width = 9 feet 
then we have,  l= 35 feet
               h= 10 feet
               W= 9  feet.  

Now volume will be = l*h*w
                   = 35*10*9
                   =3150 cubic feet